OpenStudy (anonymous):

An airplane has just enough fuel for a 5hour flight, how far can it fly roundtrip if it flies 225mph with the wind during the first leg of the trip and 180mph against the wind on the return?

OpenStudy (broskishelleh):

Set up two equations for the flight. x will represent how much time is spent on the 225mph leg of the trip, and y will represent how much time is spent on the 180mph part of the trip.

OpenStudy (broskishelleh):

225x=180y x+y=5

OpenStudy (broskishelleh):

x=5-y 225(5-y)=180y 1125-225y=180y 1125=405y y=1125/405=25/9

OpenStudy (broskishelleh):

So now what equations do I have?

OpenStudy (broskishelleh):

5-25/9=x x=20/9

OpenStudy (broskishelleh):

Plug these values into the mph they represent to get the total distance: 20/9 * 225 + 25/9 *180 = 1,000 miles